OpenStudy (anonymous):

s=p/3+7q solve for p need help to solve

OpenStudy (anonymous):

Start by substracting 7q from both sides of the equation

OpenStudy (anonymous):

7q - s =7q- 7q+p/3 7q -s = p/3 7q-s/7=p/3 q-s=p/3 *p/p q-s=3

OpenStudy (anonymous):

s = p/3 + 7q s - 7q = p/3 +7q - 7q <--- substracting 7q from both sides s - 7q = p/3 <--- so we get this 3*( s - 7q) = (p/3) * 3 <--- now we multiply both sides by 3 3s - 21q = p <--- so we get this after the multiplication by 3 p = 3s - 21q <--- we are solving for "p" , so we write the equation this way

OpenStudy (anonymous):

why did you multiple both sides by 3? let me guess so we could get rid of the p?

OpenStudy (anonymous):

We had this equation: s - 7q = p/3 and our goal was to solve for "p" . In order to isolate the "p" out of the "p/3" we had to multiply "p/3" by 3 . In order for the equation to remain valid, we must multiple everything on the other side of the equation by 3 as well. So - we multiplied both sides of the equation by 3: ( s - 7q ) * 3 = (p/3) * 3 And that's how we got: 3s - 21q = p which is the same as: p = 3s - 21q And we have solved for "p" as we were required to - now we know what "p" is equal to, in terms of the other variables "s" and "q".

OpenStudy (anonymous):

I understand how to do it know. Thanks

OpenStudy (anonymous):

welcome!